OpenStudy (anonymous):

find the following ^12sqrt(-3)^12

OpenStudy (anonymous):

\[\sqrt[12]{(-3)^{12}}\]

jimthompson5910 (jim_thompson5910):

Use the idea that \[\Large \sqrt[n]{x^n} = |x|\] where n is an even number.

OpenStudy (eyust707):

so for example: